# Yao Cheng, Qier Cui, Yunshi Lan, Xiang Li

### v1

@@ 3.4 Semantic Mind Graph Retrieval
Table 1: Human evaluation results for Review Comments Understanding (Rank-Based $\downarrow$). We highlight the best score on each dataset in bold and the runner-up score with an underline.
3.5 Hierarchical Background Graph Retrieval
We first begin the retrieval process at the theme level, aiming to extract themes related to a review comment. For example, if a reviewer questions whether the proposed method shares similarities with certain techniques in the fields of computer vision (e.g., contrastive learning in CV), we retrieve the corresponding descriptions of theme nodes to align with the reviewer’s concerns from a broader perspective.
Based on the theme-level information, we then proceed to the abstract-level retrieval, focusing on papers related to the identified themes. Note that abstracts summarize key research questions, methodologies, and conclusions, providing a concise yet comprehensive overview. Therefore, abstract-level retrieval is particularly useful for understanding review comments that compare the originality of the proposed approach with existing methods.
After retrieval across all three granularity levels, the top-$k$ relevant nodes are ranked and selected to ensure that the most pertinent evidence is used for understanding the review comments. More implementation details are provided in the Appendix B.
4 Experiments
Figure 4: Automated evaluation results for Review Comments Understanding, measured by Win-Rate ($\uparrow$).
4.1 Experimental Setup
Datasets.
We select 20 papers and their corresponding reviews from ICLR over the past five years through OpenReview [^1] https://openreview.net/. Each paper was reviewed by 3 to 8 reviewers, with a total of 82 reviewers. We divide the review comments and exclude those longer than 200 characters, which are considered relatively complete. In the end, 284 shorter comments were retained. These papers cover diverse fields such as Artificial Intelligence (AI), Machine Learning (ML), Natural Language Processing (NLP), and Computer Vision (CV). The selected papers exhibit a wide range of citation counts, including both highly cited and less-cited works, as well as recent studies and earlier publications.
Baseline Methods.
To validate the effectiveness of SEAGraph in terms of graph construction and retrieval, we compare it with the following two categories of baseline methods: (1) Direct inference methods: DirectInfer takes the review comment and the full parsed paper as input to directly reason about the understanding of each review comment. (2) RAG-based methods: RAG-naive computes the similarity between each review comment and every chunk of the paper, selecting the top-$k$ chunks to combine with the review comment as input; RAG-SMG utilizes only the construction and retrieval of the Semantic Mind Graph; RAG-HBG relies solely the construction and retrieval of the Hierarchical Background Graph.
We conduct experiments using the same open-source base model Ministral-8B-Instruct-2410222https://mistral.ai/ and the same prompt (Table 3) as input to LLMs.
Evaluation Protocol.
Since the task of reivew comments understanding lacks a definitive ground truth and exhibits significant diversity in generated content, we design two evaluation methods: human evaluation and automated evaluation. For human evaluation, we engage 40 experts from various fields, each responsible for one paper to assess the quality of the generated understanding of the review comments. For automated evaluation, given the powerful text comprehension capabilities of LLMs to play as a judge Li et al. (2024a), we employ gpt-4o-2024-11-20333https://openai.com/ as the evaluation model to provide objective judgments for the task.
4.2 Main experiments
Evaluation Metric.
There are four main assessment metrics for human and automated evaluation: (1) Relevance: Assesses the alignment between the provided evidence and the review comments. (2) Clarity: Evaluates how clearly and effectively the information is presented for ease of understanding. (3) Criticality: Examines the depth of analysis and the extent to which the feedback reflects constructive thought. (4) Novelty: Measures the inclusion of fresh insights or new evidence. Additionally, two customized human-centric metrics are tailored for evaluation: (1) Persuasiveness: Focuses on the logical reasoning and the ability of arguments to persuade. (2) Practicality: Gauges the usefulness and applicability of the information for authors. The specific meaning of six metrics is provided in Appendix A.
Figure 5: Human evaluation results for Key concerns in Reviews. ($\uparrow$).
Human Evaluation Results.
Automated Evaluation Results.
In the automated evaluation, due to the excessive total text length generated by the five methods, we employ a pairwise ranking approach for assessment. Figure 4 presents the results for review comments with a length of less than 100, while the results and analysis for lengths between 100 and 200 are provided in Appendix C. The values in the heatmap represent the win rate of the method shown on the vertical axis over the method on the horizontal axis. From the figure, it can be observed that: (1) SEAGraph consistently performs as the optimal method across the four metrics. (2) In terms of Criticality and Novelty metrics, SEAGraph significantly outperforms other methods, indicating its ability to provide more innovative evidence and conduct deeper analysis for review comments. (3) In most cases, RAG-SMG consistently ranks second. Although it performs slightly worse than RAG-naive in terms of relevance, it surpasses RAG-naive across all other metrics. This suggests that while RAG-SMG captures slightly less relevant aspects of evidence, its ability to make logical connections greatly enhances the reasoning power of LLM. These findings highlight the crucial role of modeling academic papers as semantic mind graphs to capture the paper’s underlying structure and logic. (4) Notably, for novelty, RAG-HBG performs better than all methods except SEAGraph, as it introduces multiple perspectives beyond the reviewed paper.
Overall, both human and automated evaluations yield consistent conclusions, providing strong evidence that SEAGraph is capable of generating valuable insights for interpreting review comments.
4.3 Quantitative Evaluation of Retrieval
In this section, we automate the evaluation of retrieved content based on five distinct metrics: (1) Relevance, (2) Specificity, (3) Novelty, (4) Logic, (5) Explainability. The detailed meanings are shown in Appendix A. Due to the extensive length of the retrieved content, both the content and corresponding review comment are directly input into gpt-4o for scoring. The results, as presented in the Table 2, demonstrate that RAG-SMG outperforms other methods in the majority of cases. Specifically, while RAG-SMG scores slightly lower than RAG-naive on the Relevance metric, this discrepancy may be attributed to the retrieval of chunk nodes with marginally lower semantic similarity through the logical structure of the semantic mind graph. Conversely, the external background knowledge retrieved by RAG-HBG exhibits limited relevance to the review comment, leading to consistently lower scores across various evaluation metrics. In summary, the information retrieved via the semantic mind graph demonstrates greater contextual significance and utility.
4.4 Key Concerns in Reviews
In this section, we consider the issues highlighted in the review comments as the key concerns of the paper. We first use LLMs to summarize all the reviews of each paper, identifying the concerns most frequently mentioned by the reviewers. Then, we employ six metrics consistent with those in Section 4.2 and perform a human evaluation to quantify and compare the effectiveness of  SEAGraph in understanding and explaining these concerns with other methods. The results are shown in Figure 5. It can be seen that: SEAGraph consistently achieves best scores across all evaluation metrics, particularly excelling in persuasiveness and practicality when addressing the key concerns identified in the reviews, indicating a greater alignment with human preference. RAG-SMG also shows strong performance in all metrics, as the concerns of reviewers are closely aligned with the content of the paper. In contrast, RAG-HBG has a poor performance since the supplemental background knowledge is less helpful to the key concerns. In conclusion, both SEAGraph and RAG-SMG show superior performance in this experiment, garnering higher recognition for their practicality and persuasiveness from human evaluators, further confirming the superiority of our framework in assisting authors to understand reviews.
Metric $\uparrow$ RAG-naive RAG-SMG RAG-HBG
Relevance 6.46 6.45 4.75
Specificity 5.74 5.84 4.23
Novelty 5.06 5.10 3.68
Logic 6.57 6.63 4.82
Explainability 5.57 5.62 3.94
Table 2: Automated evaluation results for retrieval.
5 Conclusion
In a nutshell, we sincerely hope that our work does empower authors to not only gain a deeper understanding of reviews feedback but also elevate the quality of their papers, ultimately expediting both the advancement of research and the efficiency of the submission process.
Limitations
Rebuttal mechanism.
The rebuttal mechanism, where authors respond to reviewers’ concerns and engage in further discussion, also plays a critical role in improving the paper Jin et al. (2024). The success of multi-agent systems in executing complex tasks presents a promising opportunity Wu et al. (2023, 2024b). In future research, we will explore simulating the rebuttal process through multi-agent communication, aiming to further bridge the understanding gap between reviewers and authors in papers and comments, thus advancing the rebuttal mechanism.
Enhancing Pipeline Stability.
Privacy Challenges in Benchmarking.
We plan to establish a standardized benchmark for the field of peer review to help standardize and advance research practices in this area. Currently, platforms such as OpenReview provide a wealth of publicly available papers and review data, offering valuable resources for studying the mechanisms and effectiveness of peer review. However, there are certain limitations to the use of these data, primarily concerning privacy protection. While the identities of authors are ostensibly anonymized, the peer review process allows senior roles such as Program Chairs and Area Chairs to access the actual identities of both authors and reviewers. This potential breach of privacy may pose challenges to the objectivity, fairness, and ethical considerations of related research.
Evaluation Metrics.
Although human evaluation and GPT-based evaluation can reflect the strengths and weaknesses of a model to some extent, they often involve significant subjectivity and lack consistency. This issue becomes particularly pronounced in open-ended generation tasks, where differences in standards and preferences among evaluators may lead to inconsistent results. Therefore, establishing a comprehensive, unified, and reproducible quantitative standard is crucial for more objective and fair assessment of model performance. Such a standard not only helps to minimize the influence of human bias but also provides more actionable feedback for subsequent model optimization and improvement.
Ethics Statement
Appendix A More details of SEAGraph

### v2

@@ 3.4 Semantic Mind Graph Retrieval
3.5 Hierarchical Background Graph Retrieval
We first begin the retrieval process at the theme level, aiming to extract themes related to a review comment. For example, if a reviewer questions whether the proposed method shares similarities with certain techniques in the fields of computer vision (e.g., contrastive learning in CV), we retrieve descriptions of related theme nodes to broadly align with the reviewer’s concerns. Based on the theme-level information, we then proceed to the abstract-level retrieval, focusing on papers related to the identified themes. Note that abstracts summarize key research questions, methodologies, and conclusions, providing a concise yet comprehensive overview. Therefore, abstract-level retrieval is particularly useful for understanding review comments that compare the originality of the proposed approach with existing methods.
After retrieval across all three granularity levels, the top-$k$ relevant nodes are ranked and selected to ensure that the most pertinent evidence is used for understanding the review comments. More implementation details are provided in the Appendix G.
4 Experiments
4.1 Experimental Setup
Datasets.
We collect a total of 1,256 review comments from ICLR submissions over the past five years via the OpenReview platform [^1] https://openreview.net/. Each comment contains no more than 200 characters, as longer reviews are typically considered sufficiently comprehensive and thus less suitable for our analysis. The associated papers span diverse areas of artificial intelligence, with citation counts ranging from highly cited works to those with minimal impact, covering both recent studies and earlier publications.
The research papers are categorized into six major areas: Natural Language Processing (NLP), Multimodal Learning (MM), Computer Vision (CV), Representation Learning (RLearn), Theory and Optimization (Theory&Opt.), and Robustness (Robust.). Figure 4 illustrates the specific proportion of each category. Further, Figure 5 presents the distribution of review lengths in our dataset. A prominent peak appears around 75 characters, indicating that a large number of short comments tend to concentrate at this length.
Figure 4: Research paper topic distribution across six key areas. NLP: Natural Language Processing; $M$: Multimodal Learning; CV: Computer Vision; RLearn: Representation Learning; Theory&Opt.: Theory and Optimization; Robust.: Robustness.
Figure 5: Distribution of Review Lengths.
Figure 6: Automated evaluation results for Review Comments Understanding, measured by Win-Rate (% $\uparrow$).
Baseline Methods.
To validate the effectiveness of SEAGraph in terms of graph construction and retrieval, we compare it with the following two categories of baseline methods: (1) Direct inference method: DirectInfer takes the review comment and the full parsed paper as input to directly reason about the understanding of each review comment. (2) RAG-based methods: RAG-naive computes the similarity between each review comment and every chunk of the paper, selecting the top-$k$ chunks to combine with the review comment as input; RAG-SMG utilizes only the construction and retrieval of the Semantic Mind Graph; RAG-HBG relies solely the construction and retrieval of the Hierarchical Background Graph.
We conduct experiments on all baseline models using the same open-source foundation models, Qwen3-8B [^2] https://qwenlm.github.io/ and Ministral-8B-Instruct-2410333https://mistral.ai/. Due to space constraints, the experimental details for Ministral are provided in Appendix A. The example of SMG and HBG construction is in Appendix E and the example of SEAGraph revealing review comments is in Appendix D.
Metric $\uparrow$ RAG-naive RAG-SMG RAG-HBG
Relevance 7.59 7.53 5.33
Specificity 7.54 7.60 5.47
Novelty 6.46 6.54 4.76
Logic 7.65 7.68 5.23
Explainability 7.06 7.09 4.61
Table 1: Automated evaluation results for retrieval.
DirectInfer 3.45 3.85 3.87 3.60 3.41 3.62 3.63
RAG-naive 3.11 2.93 2.90 3.24 3.16 2.97 3.05
RAG-SMG 2.61 2.55 2.37 2.81 2.69 2.63 2.61
RAG-HBG 3.94 3.70 3.84 3.40 3.91 3.79 3.76
SEAGraph 1.89 1.97 2.02 1.95 1.83 1.99 1.94
Table 2: Human evaluation results for Review Comments Understanding (Rank-Based $\downarrow$). We highlight the best score on each metric in bold and the runner-up score with an underline.
Evaluation Protocol.
Since the task of review comments understanding lacks a definitive ground truth and exhibits significant diversity in generated content, we design two evaluation methods: automated evaluation and human evaluation. For automated evaluation, given the powerful text comprehension capabilities of LLMs to play as a judge li2024generation, we employ gpt-4o-2024-11-20 and Qwen3-14B as the evaluation model to provide objective judgments for the task. For human evaluation, we recruit 40 experts from diverse academic background, detailed information is provided in Appendix B.
4.2 Automated Evaluation Results.
Evaluation Metric.
There are four main assessment metrics for human and automated evaluation: (1) Relevance: Assesses the alignment between the provided evidence and the review comments. (2) Clarity: Evaluates how clearly and effectively the information is presented for ease of understanding. (3) Criticality: Examines the depth of analysis and the extent to which the feedback reflects constructive thought. (4) Novelty: Measures the inclusion of fresh insights or new evidence. The specific meaning is provided in Appendix C.
Result.
In the automated evaluation, due to the excessive total text length generated by the five methods, we employ a pairwise ranking approach for assessment. Figure 6 presents the results for review comments. The values in the heatmap represent the win rate of the method shown on the vertical axis over the method on the horizontal axis. From the figure, we can see that: (1) SEAGraph consistently outperforms all baseline methods across all four evaluation metrics —Relevance, Clarity, Criticality, and Novelty— demonstrating its superior ability to understand review comments. Its win rates exceed 77% in all pairwise comparisons, with several cases reaching above 85%, indicating a strong advantage in both content alignment and critical interpretation. (2) RAG-SMG and RAG-HBG serve as moderately effective strategies, incorporating either semantic or hierarchical knowledge structures. Both methods contribute meaningful improvements, demonstrating that integrating information from distinct knowledge dimensions—internal semantics and external context—provides valuable support for interpreting review comments. They show noticeable improvements over the naive and Infer baselines, suggesting that structured input plays an important role in enhancing the model’s comprehension of review intent. (3) DirectInfer and RAG-naive methods perform poorly across all metrics, particularly in tasks requiring deeper understanding such as clarity and novelty. Their limitations stem from relying either solely on the original paper content or on unstructured retrieved evidence, both of which lack the logical organization needed to support reviewer-level reasoning. Overall, SEAGraph effectively generates insights for interpreting review comments.
Figure 7: Human evaluation results for Key concerns in Reviews ($\uparrow$).
4.3 Quantitative Evaluation of Retrieval
The evaluation of retrieved content includes five distinct metrics: (1) Relevance, (2) Specificity, (3) Novelty, (4) Logic, (5) Explainability. The detailed meanings are shown in Appendix C. Due to the extensive length of the retrieved content, both the content and corresponding review comment are directly input into gpt-4o for scoring. The results, as presented in the Table 1, demonstrate that RAG-SMG outperforms other methods in the majority of cases. Specifically, while RAG-SMG scores slightly lower than RAG-naive on the Relevance metric, this discrepancy may be attributed to the retrieval of chunk nodes with marginally lower semantic similarity through the logical structure of the semantic mind graph. Conversely, the external background knowledge retrieved by RAG-HBG exhibits limited relevance to the review comment, leading to consistently lower scores across various evaluation metrics. In summary, the information retrieved via the semantic mind graph demonstrates greater contextual significance and utility. Additionally, we include retrieval examples in Appendix F to further illustrate the advantages of the Semantic Mind Graph.
4.4 Human Evaluation Results
In the human evaluation process, experts are invited to rank the review understanding results produced by five different methods based the following metrics. We adopt the same four automatic evaluation metrics introduced in Section 4.2: Relevance, Clarity, Criticality, and Novelty. Additionally, two customized human-centric metrics are tailored for evaluation: (1) Persuasiveness: Focuses on the logical reasoning and the ability of arguments to persuade. (2) Practicality: Gauges the usefulness and applicability of the information for authors. The results are summarized in Table 2, where the scores indicate the average ranking over all samples.
We can observe that: (1) SEAGraph achieves the highest performance across all metrics, underscoring the effectiveness of our framework. (2) In terms of average rankings, RAG-SMG performs better than RAG-naive, which in turn outperforms DirectInfer. These findings strongly support our motivation that logically retrieved content significantly improves the ability of LLMs to understand review comments. (3) Although RAG-HBG performs poorly—mainly because its retrieved content consists only of background knowledge without incorporating the internal knowledge of the paper—SEAGraph enhances the LLM’s ability to understand the paper by integrating external background knowledge on top of RAG-SMG, enabling it to extract more meaningful information for reasoning. Consequently, constructing the semantic mind graph and the hierarchical background graph can provide valuable support for understanding paper reviews from different perspectives.
5 Analysis of Human Evaluation Consistency
To assess the reliability and effectiveness of human evaluation in our study, we analyze the NDCG@5 scores across Qwen3-8B. NDCG@5 is computed based on independent human judgments for the model’s generated outputs. In the context of human evaluation, this metric quantifies how well annotators can distinguish and rank high-quality model outputs from lower-quality ones within the model’s generated responses. By computing NDCG@5 scores across multiple evaluation criteria, we aim to evaluate both the overall reliability and the sensitivity of human judgments in capturing meaningful differences in response quality.
The results are summarized in Table 3 and reflect average NDCG@5 scores across six evaluation criteria: Relevance, Clarity, Critical Insight, Novelty, Persuasiveness, and Practicality. As shown in the table, Qwen3-8B achieves relatively high NDCG@5 scores across all evaluation criteria, indicating that human annotators were able to consistently rank the quality of its generated responses. These results suggest strong reliability and agreement among annotators, likely because the model produces stable and coherent outputs that reduce ambiguity during evaluation.
Metric Qwen3-8B
Relevance 0.9105
Clarity 0.9066
Critical Insight 0.8991
Novelty 0.8991
Persuasiveness 0.8874
Practicality 0.8995
Table 3: NDCG@5 of Qwen3 based on human judgments.
5.1 Key Concerns in Reviews
In this section, we consider the issues highlighted in the review comments as the key concerns of the paper. We first use LLMs to summarize all the reviews of each paper, identifying the concerns most frequently mentioned by the reviewers. Then, we employ six metrics consistent with those in Section 4.4 and perform a human evaluation to quantify and compare the effectiveness of  SEAGraph in understanding and explaining these concerns with other methods. The results, shown in Figure 7, indicate that SEAGraph consistently achieves best scores across all evaluation metrics, particularly excelling in persuasiveness and practicality, demonstrating strong alignment with human preferences. RAG-SMG also shows strong performance in all metrics, as the concerns of reviewers are closely aligned with the content of the paper. In contrast, RAG-HBG has a poor performance since the supplemental background knowledge is less helpful to the key concerns. Overall, both SEAGraph and RAG-SMG show superior performance, garnering higher recognition for their practicality and persuasiveness from human evaluators, further confirming the superiority of our framework in assisting authors to understand reviews.
6 Conclusion
In a nutshell, we sincerely hope that our work does empower authors to not only gain a deeper understanding of reviews feedback but also elevate the quality of their papers, ultimately expediting both the advancement of research and the efficiency of the submission process. We hope this work can inspire further efforts toward making peer review more accessible, informative, and impactful.
Limitations
Rebuttal mechanism.
The rebuttal mechanism, where authors respond to reviewers’ concerns and engage in further discussion, also plays a critical role in improving the paper jin2024agentreview. The success of multi-agent systems in executing complex tasks presents a promising opportunity autogen; oscopilot. In future research, we will explore simulating the rebuttal process through multi-agent communication, aiming to further bridge the understanding gap between reviewers and authors in papers and comments, thus advancing the rebuttal mechanism.
Enhancing Pipeline Stability.
More granular Evaluation.
Although human evaluation and GPT-based evaluation can reflect the strengths and weaknesses of a model to some extent, they often involve significant subjectivity and lack consistency. This issue becomes particularly pronounced in open-ended generation tasks, where differences in standards and preferences among evaluators may lead to inconsistent results. Therefore, establishing a comprehensive, unified, and reproducible quantitative standard is crucial for more objective and fair assessment of model performance. Such a standard not only helps to minimize the influence of human bias but also provides more actionable feedback for subsequent model optimization and improvement. In addition, a comprehensive evaluation requires more data and human involvement, which brings with it significant costs.
Challenges in Benchmarking.
We plan to establish a standardized benchmark for the field of peer review in the future to help standardize and advance research practices in this area. Currently, platforms such as OpenReview provide a wealth of publicly available papers and review data, offering valuable resources for studying the mechanisms and effectiveness of peer review. However, there are certain limitations to the use of these data. On one hand, there is the issue of data incompleteness. Typically, an article will have three or four reviewers, or even more, and extracting and merging the key information from these reviews poses a significant challenge. Meanwhile, for review comments understanding task, the ground truth needs to explicitly capture the reasoning behind each reviewer’s comments and form a complete logical chain. This process not only requires the flexible use of LLMs for assistance but also demands the deep involvement of experts, which brings with it significant and potentially immeasurable costs. On the other hand, there are concerns regarding privacy protection. Although the identities of authors are ostensibly anonymized, the peer review process allows senior roles such as Program Chairs and Area Chairs to access the actual identities of both authors and reviewers. This potential breach of privacy may pose challenges to the objectivity, fairness, and ethical considerations of related research.
Ethics Statement
Acknowledgments
This work is supported by National Natural Science Foundation of China (No. 62202172).
Appendix A Performance with Ministral-8B
In this section, we evaluate a subset of 284 review comments sampled from the original dataset using a different large language model, Ministral-8B-Instruct-2410444https://mistral.ai/. The same prompt from Section 4.2 is adopted to ensure consistency. Evaluation metrics remain aligned with those defined in Sections 4.2 and 4.4. For automated evaluation, we utilize gpt-4o-2024-11-20555https://openai.com/ as the evaluation model to generate objective and standardized judgments across all dimensions.
Figure 8: Automated evaluation results for Review Comments Understanding, measured by Win-Rate (% $\uparrow$), based on comments with length <100<100.
Figure 9: Automated evaluation results for Review Comments Understanding, measured by Win-Rate (% $\uparrow$), based on comments with length in the range of $(100,200)$.
A.1 Automated Evaluation Results.
In the automated evaluation, due to the excessive total text length generated by the five methods, we employ a pairwise ranking approach for assessment. Figure 8 presents the results for review comments with a length of less than 100, while the results. The values in the heatmap represent the win rate of the method shown on the vertical axis over the method on the horizontal axis. From the figure, it can be observed that: (1) SEAGraph consistently performs as the optimal method across the four metrics. (2) In terms of Criticality and Novelty metrics, SEAGraph significantly outperforms other methods, indicating its ability to provide more innovative evidence and conduct deeper analysis for review comments. (3) In most cases, RAG-SMG consistently ranks second. Although it performs slightly worse than RAG-naive in terms of relevance, it surpasses RAG-naive across all other metrics. This suggests that while RAG-SMG captures slightly less relevant aspects of evidence, its ability to make logical connections greatly enhances the reasoning power of LLM. These findings highlight the crucial role of modeling academic papers as semantic mind graphs to capture the paper’s underlying structure and logic. (4) Notably, for novelty, RAG-HBG performs better than all methods except SEAGraph, as it introduces multiple perspectives beyond the reviewed paper.
Figure 9 shows the results for samples of all lengths. From the figure, we can see that SEAGraph outperforms in most cases, only slightly falling short on the Relevance metric when compared to Infer. On the other hand, RAG-SMG lags slightly behind RAG-naive in terms of relevance, likely because longer review comments are already sufficiently detailed, leading to a minor disadvantage in argument relevance. However, both SEAGraph and RAG-SMG demonstrate superior performance on other metrics, proving that the evidence they provide are more effective and better support reasoning.
A.2 Human Evaluation Results.
Overall, both automated and human evaluations yield consistent conclusions, providing strong evidence that SEAGraph is capable of generating valuable insights for interpreting review comments.
A.3 Analysis of Human Evaluation Consistency
To assess the reliability and effectiveness of human evaluation in our study, we conduct a detailed analysis of NDCG@5 scores for Ministral-8B. NDCG@5 is computed based on independent human judgments for the model’s generated outputs. This metric quantifies how well annotators can distinguish and rank high-quality responses from lower-quality ones within the model’s output set. By examining NDCG@5 scores across multiple evaluation criteria, we aim to understand the overall reliability of human judgments and their sensitivity to meaningful variations in response quality.
The results, summarized in Table 4, show the average NDCG@5 scores of Ministral-8B across six evaluation criteria: Relevance, Clarity, Critical Insight, Novelty, Persuasiveness, and Practicality. The consistently high scores across all dimensions suggest that human annotators were able to reliably differentiate response quality within the model’s generated outputs. This indicates strong agreement among annotators and provides evidence that the evaluation protocol captures stable and interpretable human judgments. Such consistency reinforces the credibility of our human evaluation setup and supports the reliability of conclusions drawn from it.
Metric Ministral-8B
Relevance 0.8761
Clarity 0.8600
Critical Insight 0.8785
Novelty 0.8616
Persuasiveness 0.8664
Practicality 0.8675
Table 4: NDCG@5 of Ministral based on human evaluation.
Appendix B Experts information for human evaluation
For human evaluation, we recruit 40 experts, including master’s and doctoral students from diverse academic backgrounds. Each review comment is independently evaluated by two experts to assess the quality of understanding. All participants with prior experience in publishing peer-reviewed papers or serving as academic conference reviewers were compensated at a rate of \$10 per hour.
@@ Appendix B Experts information for human evaluation
Table 5: Human evaluation results for Review Comments Understanding (Rank-Based $\downarrow$). We highlight the best score on each metric in bold and the runner-up score with an underline.
Appendix C More Details of SEAGraph